OpenStudy (anonymous):

Perform the subtraction and then use fundamental identities to simplify the following. 1/(sec x +) - 1/ (sec x -1) AND tan x - sec 2x/tanx

OpenStudy (anonymous):

\[1\div \sec x + 1 - 1\div \sec x-1\] I screwed up the first problem.

OpenStudy (cwrw238):

1 1 ----- - ----- secx + 1 secx - 1 secx - 1 - (secx + 1) = ------------------- sec^2 x - 1 -2 = ------------- tan^2 x

OpenStudy (cwrw238):

is second part tan x - sec 2x ------- tan x ?

OpenStudy (anonymous):

No it's not. secant squared x.

OpenStudy (cwrw238):

tan x - sec^2 x ------ tan x tan^2 x - sec^2 = -------------- tan x = tan^2 x - ( 1 + tan^2 x) --------------- tan x = -1 ---- = - cot x tanx

OpenStudy (anonymous):

Now how would I do that? Is there a chart or something with all these things that says so and so is equal to something?

OpenStudy (cwrw238):

these are trigonometrical identities do google search - they should be on wikipoedia also) examples are sec^2x = 1 + tan^2x (the one i used above) sin^2x + cos^2x = 1 cosec^2 x = 1 + cot^2 x tan x = sin x / cos x - and quite a few more

OpenStudy (anonymous):

Damn man I gotta memorize all this junk alright I'll do my best.

OpenStudy (cwrw238):

yea - i'm afraid so